a) A 150Ω resistor is connected in series to a1kΩ resistor and to a battery of 127V. Find the effective resistance and current in the circuit.

b) A 10Ω resistor is connected in parallel to 32Ω resistor and a voltage source of 24V. Find the effective resistance and total current provided in the circuit. Determine the current in each of the resistors.



1
Expert's answer
2021-11-14T17:49:30-0500

Part (a)


For series arrangement of reistors

"R_{eff}=R_1+R_2"

"=150+1000=1150\Omega"


Current in circuit

"\Iota=\frac{e.m.f.\>of voltage\> source}{total\> resistance}"


"=\frac{127}{1150}=0.11A"


Part B

Effective resistance for two resistors in parallel


"R_{eff}=\frac{R_1R_2}{R_1+R_2}=\frac{10×32}{10+32}=7.62\Omega"


Total current


"\Iota_T=\frac{e.m.f\>of\>Voltage \>source}{effective\,resistance}"


"=\frac{24}{7.62}=3.15A"


Current through a resistor in parallel arrangement


"\Iota_R=\frac{pd \>across\> the\> parallel \>arrangement}{Resistance \>of \>the \>resistor}"


"\Iota_{10\Omega}=\frac{24}{10}=2.4A"


"\Iota_{32\Omega}=\frac{24}{32}=0.75A"